OpenStudy (anonymous):

The table below shows the ticket rates for whale watching trips offered by Dolphin Tricks & Tours. Age Ticket - Price Under three years - free Three to twelve years - $35 Over twelve years $46 On a certain day, the company took 88 people on whale watching trips. There were 44 children aged twelve and under, of which some children were under three years. If x represents the number of children under three years, which equation can be used to find the value of x, where C represents the total amount of mo

OpenStudy (anonymous):

not enough information

OpenStudy (anonymous):

From the amount of ...... money collected from tickets that day? (44 + x)35 + (88 − (44 + x)) ∙ 46 = C (88 − x)35 + (88 − 44) ∙ 46 = C (44 − x)35 + (88 − 44) ∙ 46 = C (88 + x)35 + (88 − (44 + x)) ∙ 46 = C